JL.51 Carls Custom Cans produces small containers which are purchased by candy and snack food producers. The production faciPlease show your work so I can understand this portion, Thank you!
Answer #1

1st question

Economic production quantity(EPQ) = Sqrt(((2*D*Co)/Ch)*(p/(p-d)) where p = production rate = 120, d = daily demand = Annual demand/no of working days = 13800/260 = 53.07692308,D = annual demand = 13800

,Co = ordering cost = 43.31, Ch = holding cost = 2.3


Economic production quantity = Optimal production lot size = SQRT(((2*13800*43.31)/2.3)*(120/(120-53.07692308))) = 965.3557501 = 965 (Rounded to nearest whole number)

2nd question

Number of production runs = Annual demand/Optimal production lot size = 13800/965.3557501=14.29524815 = 15 (Rounded to next whole number)

3rd question

Total cost with EPQ =(1/2)(EPQ/p)(p-d)Ch+(D/EPQ)Co = (1/2)(965.3557501/120)*(120-53.07692308)*2.3+(13800/965.3557501)*43.31 = 1238.254395

To have order quantity 800 as optimal production quantity, total cost with quantity 800 needs to be 1238.254395

Let, reduced order cost with this = S

So, 2*(Annual demand/production quantity)*S = 1238.254395

or, 2*(13800/800)*s = 1238.254395

or, s = 1238.254395/34.5 = 35.89143174

Required set up cost = 35.89 (Rounded to 2 decimal places)

4th question

As Carl pays $18 per hour for setup labor, time it should take to set up this production line = (35.89/18) = 1.993888889 hour = 1.993888889*60 minutes = 119.6333333 minutes = 120 minutes (rounded to nearest whole number)